# Modeling User Preferences as Distributions for Optimal Transport-based Cross-domain Recommendation under Non-overlapping Settings

### v2

@@ Abstract
Experiments on Amazon Review datasets demonstrate that DUP-OT mitigates domain discrepancy and significantly outperforms state-of-the-art baselines under strictly non-overlapping training settings, with user correspondence revealed only for inference-time evaluation. Code is available at https://github.com/XiaoZY2000/DUP-OT-organized.
Keywords Cross-Domain Recommendation $\cdot$ Non-overlapping Users $\cdot$ Gaussian Mixture Model (GMM) $\cdot$ Optimal Transport $\cdot$ Cold-start Problem
@@ 1 Introduction
With the rapid growth of online information, recommender systems have become essential for helping users navigate massive content R_Lü_2012. However, conventional single-domain recommenders suffer from cold-start and data sparsity issues T_He_2024; L_Liu_2024; A_Zang_2023. Cross-domain recommendation (CDR) alleviates these problems by transferring knowledge from a dense source domain to a sparse target domain A_Zhu_2018. Yet most existing CDR methods rely on overlapping users or items in the training stage to bridge domains and typically represent user preferences as discrete vectors, which limits their ability to capture multi-aspect preferences.
In this work, we study the non-overlapping CDR scenario—where no users or items are shared across domains—and propose to represent user preferences using Gaussian Mixture Models (GMMs). This enables fine-grained, multi-aspect preference modeling, but also introduces two challenges: (1) domain discrepancy becomes harder to address without overlapping entities D_Guo_2024; C_Liu_2023, and (2) GMM-based preference modeling significantly increases computation and is not directly compatible with standard optimal transport formulations.
Recent CDR approaches attempt to reduce domain discrepancy using contrastive learning, adversarial training, or distribution alignment L_Liu_2024; R_Wang_2020; C_Perera_2019. Other studies explore distributional representations but either adopt simple Gaussian models or still rely on overlapping entities M_Zhu_2024. Thus, how to build expressive GMM-based preference representations while enabling cross-domain alignment in a strictly non-overlapping setting remains an open problem.
@@
  • We validate DUP-OT on Amazon Review datasets, showing consistent performance gains over baselines.
2 Related Works
Cross-Domain Recommendation. CDR aims to transfer knowledge from a rich source domain to a sparse target domain. Early approaches rely on overlapped users or items through representation alignment D_Li_2020, adversarial learning R_Li_2022, and graph-based propagation D_Guo_2021. However, such overlaps are rarely available in practice.
Non-overlapping CDR methods typically learn shared latent spaces using embedding alignment L_Wang_2021, generative modeling T_Salah_2021, or distribution matching M_Zhu_2024. While recent work explores Gaussian-based representations, they either use simple single-Gaussian models or still rely on partial overlaps. Modeling user preferences with expressive GMMs under fully non-overlapping settings remains underexplored.
Optimal Transport in CDR. OT provides a principled tool for aligning distributions across domains. It has been applied to match user/item embeddings G_Li_2022 or latent distributions L_Liu_2024. However, prior work does not integrate OT with GMM-based user preference modeling, nor address the computational challenges in fully non-overlapping scenarios.
@@ 3 Preliminary
Definitions of notations. We consider two domains: the source domain $s$ and the target domain $t$. For each domain $x\in\{s,t\}$, we denote the user set and item set as $\mathcal{U}_{x}$ and $\mathcal{V}_{x}$. Although users or items may exist in both domains in reality, in our CDR setting we impose that the model does not use any overlapping-user or overlapping-item information during training. Therefore, the training process treats the two domains as non-overlapping, even if overlaps may exist in the raw data.
@@
Target-domain dataset split. The target-domain interactions are divided into a training set $\mathcal{D}_{t}^{\mathrm{train}}$, a validation set $\mathcal{D}_{t}^{\mathrm{valid}}$, and a test set $\mathcal{D}_{t}^{\mathrm{test}}$. During training, the model only uses $\mathcal{D}_{s}^{\mathrm{train}}$ and $\mathcal{D}_{t}^{\mathrm{train}}$, and any user/item overlaps between the two domains are explicitly ignored. However, validation and test sets may contain users or items that also appear in the source domain, since they are used only for evaluation and not for model optimization.
Practical motivation. This setting corresponds to real-world scenarios where user or item identities might be shared across platforms, but privacy constraints or system limitations prevent such cross-domain links from being used during training. For example, users may later appear in the target domain after the model has been trained, or items may be synchronized across platforms only at deployment time. Thus, overlaps can exist for evaluation but are not exploited for learning, faithfully reflecting practical cross-domain recommendation conditions.
Task Formulation. We aim to address the cold-start problem in the target domain by transferring knowledge from the source domain under a non-overlapping CDR setting. Specifically, the goal is to leverage source-domain data to enhance the rating prediction performance in the target domain.
@@ 4.2 Shared Preprocessing Stage
To capture semantic similarity across domains, we derive user and item embeddings from review text. For each user and item, all associated reviews are encoded using a shared pre-trained sentence encoder, and we apply a time-aware weighting scheme so that recent reviews contribute more to the final representation.
@@ 4.3 User GMM Weights Learning Stage
4.4 Cross Domain Rating Prediction Stage
@@
OT-based alignment of Gaussian components. Since all users within a domain share the same set of Gaussian components, we apply OT only between the two domains’ GMM component sets, avoiding instance-level transport. The OT problem is formulated as Eq. 4.
@@
Three test-time cases. For each user in the target-domain test set, we construct a final user distribution in the target domain depending on the interaction availability:
@@ 5 Experiments and Results
In this section, we conduct experiments on our proposed model and several baseline models and compare the results to validate our model’s effectiveness. Specifically, the research questions we focus on are as follows:
@@
RQ2. How does the performance of our proposed method compare with that of existing CDR models?
To answer RQ1, we conduct an ablation study comparing target-domain performance with and without source-domain information. To address RQ2, we evaluate our target-domain model (using distribution-based user preferences but without source-domain data) against single-domain baselines. To address RQ3, we compare our proposed DUP-OT model with representative cross-domain recommendation baselines.
@@ 5.1 Experimental Setup
In this section, we introduce some details about our experiments’ setup, including datasets, model setting, baseline models, metrics.
@@
TDAR 1.5688 1.0904 1.5677 1.0896 1.5688 1.0902
DUP-OT (w/o source) 1.3699 0.9463 1.3494 0.8806 1.3716 0.9481
DUP-OT (w/ source) 1.2919 0.8965 1.2907 0.8774 1.3376 1.0032
@@ 5.1.1 Datasets
We use the Amazon Review 5-core datasets J_Ni_2019: Digital Music, Movies and TV, Video Games, and Electronics. Electronics is used as the target domain, and the other three as source domains. Target-domain data are split into train/validation/test (8:1:1) chronologically. To keep temporal consistency, we filter out source-domain interactions occurring later than the earliest validation timestamp in the target domain. The filtered source data and target-domain training set are used to train DUP-OT, and evaluation is performed on the target-domain test set.
@@ 5.1.3 Training Settings
We train all models using the Adam optimizer with a learning rate of $3\times 10^{-4}$ and a batch size of 256, with early stopping based on target-domain validation loss. The sentence encoder is frozen, while the shared autoencoder is jointly trained across domains to obtain unified low-dimensional embeddings. For each domain, we fit item-level Gaussian Mixture Models using the BayesianGaussianMixture estimator and train domain-specific MLPs for predicting user GMM weights and ratings. Users appearing only in validation or test sets are treated as cold-start users. All models are implemented in PyTorch and trained on a single NVIDIA GPU. To ensure robustness, all experiments are run with five random seeds $\{7,11,17,33,2025\}$, and we report the mean of all metrics.
@@ 5.1.4 Baselines
Single Domain Models:
@@
Non-overlapping Cross Domain Models:
@@
For single domain RS models, we utilize RecBole recbole[1.0]; recbole[2.0]; recbole[1.2.1] in our experiments to implement them. And for TDAR, we just manually implement them.
@@ 5.1.6 Metrics
To evaluate the performance of our proposed DUP-OT and baseline models, we set all models’ targets as rating prediction on the target domain’s test dataset. And then we use RMSE and RAE as metrics to evaluate models’ performance. The results are shown in Table 1.
@@ 5.2 Discussion of Results
Answer to RQ2 (Effect of distribution-based user modeling). Comparing DUP-OT(w/o source) with single-domain recommenders (LightGCN and NeuMF), DUP-OT(w/o source) achieves substantially better performance in the Electronics-only setting, despite using no source-domain data. This indicates that modeling user preferences as distributions rather than discrete vectors yields more expressive representations and improves rating prediction accuracy in the target domain.
Answer to RQ3 (Comparison with cross-domain baselines). Relative to TDAR, which is a representative CDR method, DUP-OT(w source) achieves consistently lower RMSE and MAE across all dataset combinations. This demonstrates that combining GMM-based preference modeling with optimal transport provides more effective cross-domain alignment than conventional approaches, validating the overall superiority of our framework.
@@ 6 Conclusion
We proposed DUP-OT, a distribution-based framework for non-overlapping cross-domain recommendation. By modeling user preferences as GMMs and aligning domains via Optimal Transport, DUP-OT enables expressive preference representation and effective cross-domain transfer without relying on shared users or items during the training stage. Experiments on Amazon Review datasets show that DUP-OT improves target-domain rating prediction, outperforms single-domain models even without source data, and achieves superior performance over a representative non-overlapping CDR baseline.

### v3

# arXiv:2508.16210v3[cs.IR] 28 Feb 2026

@@ Abstract
Experiments on Amazon Review datasets show that DUP-OT outperforms single-domain baselines even without source-domain data, and achieves lower RMSE than the cross-domain baseline TDAR under strictly non-overlapping training settings, demonstrating its effectiveness in reducing large prediction errors for cold-start users. The implementation is available at https://github.com/XiaoZY2000/dup-ot.
@@ 1 Introduction
With the rapid growth of online information, recommender systems have become essential for helping users navigate massive content [R_Lü_2012]. However, conventional single-domain recommenders suffer from cold-start and data sparsity issues [3, 10, 18]. Cross-domain recommendation (CDR) alleviates these problems by transferring knowledge from a dense source domain to a sparse target domain [21]. Yet most existing CDR methods rely on overlapping users or items in the training stage to bridge domains and typically represent user preferences as discrete vectors, which limits their ability to capture multi-aspect preferences.
In this work, we study the non-overlapping CDR scenario—where no users or items are shared across domains—and propose to represent user preferences using Gaussian Mixture Models (GMMs). This enables fine-grained, multi-aspect preference modeling, but also introduces two challenges: (1) domain discrepancy becomes harder to address without overlapping entities [1, 9], and (2) GMM-based preference modeling significantly increases computation and is not directly compatible with standard optimal transport formulations.
Recent CDR approaches attempt to reduce domain discrepancy using contrastive learning, adversarial training, or distribution alignment [10, 14, 12]. Other studies explore distributional representations but either adopt simple Gaussian models or still rely on overlapping entities [22]. Thus, how to build expressive GMM-based preference representations while enabling cross-domain alignment in a strictly non-overlapping setting remains an open problem.
@@
  • We validate DUP-OT on Amazon Review datasets, demonstrating improved RMSE over both single-domain and cross-domain baselines, with particular effectiveness in reducing large prediction errors for cold-start users.
2 Related Work
2.0.1 Cross-Domain Recommendation.
CDR aims to transfer knowledge from a rich source domain to a sparse target domain. Early approaches rely on overlapped users or items through representation alignment [7], adversarial learning [6], and graph-based propagation [2]. However, such overlaps are rarely available in practice.
Non-overlapping CDR methods typically learn shared latent spaces using embedding alignment [15], generative modeling [13], or distribution matching [22]. While recent work explores Gaussian-based representations, they either use simple single-Gaussian models or still rely on partial overlaps. Modeling user preferences with expressive GMMs under fully non-overlapping settings remains underexplored.
2.0.2 Optimal Transport in CDR.
OT provides a principled tool for aligning distributions across domains. It has been applied to match user/item embeddings [8] or latent distributions [10]. However, prior work does not integrate OT with GMM-based user preference modeling, nor address the computational challenges in fully non-overlapping scenarios.
@@ 3 Preliminary
3.0.1 Definitions of Notations.
We consider two domains: the source domain $s$ and the target domain $t$. For each domain $x\in\{s,t\}$, we denote the user set and item set as $\mathcal{U}_{x}$ and $\mathcal{V}_{x}$. Although users or items may exist in both domains in reality, in our CDR setting we impose that the model does not use any overlapping-user or overlapping-item information during training. Therefore, the training process treats the two domains as non-overlapping, even if overlaps may exist in the raw data.
@@
3.0.2 Practical Motivation.
This setting corresponds to real-world scenarios where user or item identities might be shared across platforms, but privacy constraints or system limitations prevent such cross-domain links from being used during training. For example, users may later appear in the target domain after the model has been trained, or items may be synchronized across platforms only at deployment time. Thus, overlaps can exist for evaluation but are not exploited for learning, faithfully reflecting practical cross-domain recommendation conditions.
3.0.3 Task Formulation.
We aim to address the cold-start problem in the target domain by transferring knowledge from the source domain under a non-overlapping CDR setting. Specifically, the goal is to leverage source-domain data to enhance the rating prediction performance in the target domain.
@@ 4.2 Shared Preprocessing Stage
To capture semantic similarity across domains, we derive user and item embeddings from review text. For each user and item, all associated reviews are encoded using a shared pre-trained sentence encoder, and we apply a time-aware weighting scheme so that recent reviews contribute more to the final representation. Specifically, more recent interactions are assigned higher weights when aggregating review embeddings.
@@ 4.3 User GMM Weights Learning Stage
4.4 Cross-Domain Rating Prediction Stage
@@
4.4.1 OT-Based Alignment of Gaussian Components.
Since all users within a domain share the same set of Gaussian components, we apply OT only between the two domains’ GMM component sets, avoiding instance-level transport. The OT problem is formulated as Eq. 4.
@@
4.4.2 Three Test-Time Cases.
For each user in the target-domain test set, we construct a final user distribution in the target domain depending on the interaction availability:
@@
Since all users within a domain share the same set of Gaussian components and only learn personalized mixture weights rather than full GMM parameters, the per-user computation is comparable to standard MLP-based models. Moreover, OT alignment operates only on the $m\times n$ component-level cost matrix (where $m,n\ll|\mathcal{U}|$), so the transport step adds negligible overhead compared to user-level MLP training.
@@ 5 Experiments and Results
We evaluate DUP-OT against several baselines to answer the following research questions:
@@
RQ3. How does the performance of our proposed method compare with that of existing CDR models?
To answer RQ1, we conduct an ablation study comparing target-domain performance with and without source-domain information. To address RQ2, we evaluate our target-domain model (using distribution-based user preferences but without source-domain data) against single-domain baselines. To address RQ3, we compare our proposed DUP-OT model with a representative cross-domain recommendation baseline: TDAR.
@@ 5.1 Experimental Setup
TDAR 1.2512 0.8349 1.2550 0.8378 1.2484 0.8240
DUP-OT (w/o source) 1.4632 1.0658 1.4055 0.9734 1.4407 1.0354
DUP-OT (w/ source) 1.2283 0.9315 1.2499 1.0332 1.2385 0.9815
@@ 5.1.1 Datasets
We use the Amazon Review 5-core datasets [11]: Digital Music, Movies and TV, Video Games, and Electronics. Electronics is used as the target domain, and the other three as source domains. Target-domain data are split into train/validation/test (8:1:1) chronologically. To keep temporal consistency, we filter out source-domain interactions occurring later than the earliest validation timestamp in the target domain. The filtered source data and target-domain training set are used to train DUP-OT, and evaluation is performed on the target-domain test set.
@@ 5.1.3 Training Settings
We train all models using the Adam optimizer with a learning rate of $3\times 10^{-4}$ and a batch size of 256, with early stopping based on target-domain validation loss. The sentence encoder is frozen, while the shared autoencoder is jointly trained across domains to obtain unified low-dimensional embeddings. For each domain, we fit item-level Gaussian Mixture Models using the BayesianGaussianMixture estimator and train domain-specific MLPs for predicting user GMM weights and ratings. Users appearing only in validation or test sets are treated as cold-start users. All models are implemented in PyTorch and trained on a single NVIDIA GPU.
@@ 5.1.4 Baselines
Single-Domain Models:
@@
Non-Overlapping Cross-Domain Models:
@@
For single domain RS models, we utilize RecBole [20, 19, 16] in our experiments to implement them. For TDAR, we implemented the method ourselves based on the original paper.
@@ 5.1.6 Metrics
To evaluate the performance of our proposed DUP-OT and baseline models, we set all models’ targets as rating prediction on the target domain’s test dataset. And then we use RMSE and MAE as metrics to evaluate models’ performance. The results are shown in Table 1.
@@ 5.2 Discussion of Results
5.2.2 Answer to RQ2 (Effect of Distribution-Based User Modeling).
Comparing DUP-OT(w/o source) with single-domain recommenders (LightGCN and NeuMF), DUP-OT(w/o source) achieves substantially better performance in the Electronics-only setting, despite using no source-domain data. This indicates that modeling user preferences as distributions rather than discrete vectors yields more expressive representations and improves rating prediction accuracy in the target domain.
5.2.3 Answer to RQ3 (Comparison with the Cross-Domain Baseline).
Compared with TDAR, DUP-OT (w/ source) achieves lower RMSE but higher MAE across all dataset combinations. RMSE is more sensitive to large errors, while MAE treats all errors equally. The lower RMSE indicates that DUP-OT effectively avoids extreme mispredictions, which are common for cold-start users with little interaction history. This can be explained by our GMM-based preference modeling: distributional representations provide more robust estimates for data-sparse users than point vectors, reducing the risk of severe prediction failures. On the other hand, TDAR’s joint cross-domain optimization tends to produce more evenly spread errors, leading to lower MAE but less protection against worst-case predictions. This trade-off aligns well with the cold-start scenario that DUP-OT targets.
@@ 6 Conclusion
We proposed DUP-OT, a distribution-based framework for non-overlapping cross-domain recommendation. By modeling user preferences as GMMs and aligning domains via Optimal Transport, DUP-OT enables expressive preference representation and effective cross-domain transfer without relying on shared users or items during training. Experiments on Amazon Review datasets show that DUP-OT outperforms single-domain baselines even without source-domain data, and achieves lower RMSE than the cross-domain baseline TDAR, indicating stronger robustness against large prediction errors in cold-start scenarios.
A current limitation is that the linear fusion of transferred and target-domain distributions may not be optimal for all users. Future work includes exploring adaptive fusion strategies, extending DUP-OT to implicit feedback settings, and incorporating structure-aware transport costs such as Gromov-Wasserstein distance for better cross-domain alignment.
Disclosure of Interests.
The authors have no competing interests to declare.